### v1

@@ Abstract
Large, high-purity single-crystals of hexagonal BN (h-BN) are essential for exploiting its many desirable and interesting properties. Here, we demonstrate via X-ray tomography, X-ray diffraction and scanning electron microscopy that h-BN crystals can be grown by traveling-solvent floating-zone (TSFZ). The diameters of grown boules range from 3 – 5 mm with lengths from 2 – 7 mm. Tomography indicates variable grain sizes within the boules, with the largest having areas of $\approx$ 1 mm $\times$ 2 mm and thickness $\approx$ 0.5 mm. Although the boules contain macroscale flux inclusions, the h-BN lattice itself is of high quality for samples grown under optimized conditions. The currently optimized growth procedure employs an Fe flux, moderate N2 pressure ($P_{N2}\approx$ 6 bar), and a growth rate of 0.1 mm/h. Raman spectroscopy for an optimized sample gives an average linewidth of 7.7(2) cm-1 for the E2g intralayer mode at 1365.46(4) cm-1 and 1.0(1) cm-1 for the E2g interlayer shear mode at 51.78(9) cm-1. The corresponding photoluminescence spectrum shows sharp phonon-assisted free exciton peaks and minimal signal in the energy range corresponding to carbon-related defects ($E$ = 3.9 – 4.1 eV). Our work demonstrates the viability of growing h-BN by the TSFZ technique, thereby opening a new route towards larger, high-quality crystals and advancing the state of h-BN related research.
@@ 1 Introduction
The utility of hexagonal BN (h-BN) has been recognized since at least the 1950’s, particularly for high-temperature applications, where its chemical inertness, thermodynamic stability and high electrical resistivity are extremely useful. The arrival of graphene [1, 2] subsequently stimulated additional interest in h-BN as an ideal insulating substrate for device applications [3]. h-BN is an excellent substrate for graphene-based devices due to the combination of a small lattice mismatch and atomic-level flatness (i.e. no dangling bonds) with good dielectric [1, 2, 4] and thermal transport [5, 6] properties. The chemical inertness and oxidation resistance of h-BN [7] also make it valuable for capping and encapsulation purposes [8]. Furthermore, h-BN can also take an active role, as in graphene/h-BN moiré superlattices [9, 10, 11, 12]. The same properties which make h-BN ideal for graphene also make it a useful platform for device creation using other two-dimensional (2D) materials.
More recently, h-BN is under study for a host of additional applications, including, among others, photonics, flexible electronics, non-linear optics and quantum information science and sensing technologies [13, 14, 15, 16]. A number of these applications exploit the intrinsic properties of h-BN on their own. In particular, as a naturally hyperbolic  [^1] Here, hyperbolic indicates that the relative permittivity changes sign as well as magnitude along different crystallographic directions. birefringent material in the mid-infrared range h-BN is well-suited to many photonics applications [17], such as confinement and manipulation of phonon polaritons [18]. Relatedly, point defects within h-BN have also begun to generate significant interest as single photon emitters [19, 20, 21], which have the potential to serve as an optical platform for quantum information [22] and as a spin defect for use in atomic-resolution quantum sensing [23, 24, 25].
@@
Accordingly, the primary source of high-quality h-BN single crystals for the most demanding experiments has principally been the crucible-based flux (CBF) method developed by Taniguchi and Watanabe [38], which utilizes a Ba3BN2 flux in Mo crucibles at high temperature and GPa-scale pressures. However, subsequent work has demonstrated that single-crystal h-BN samples of comparable quality and size can be made with various transition metal based fluxes (typical alloys are Fe-Cr or Ni-Cr) [39, 40, 41, 42, 43, 44, 45]. These crystals have been utilized successfully in, among others, a number of interesting photonics experiments [46, 47]. Despite these accomplishments, it is clear that further improvements are needed to enable the most ambitious technological proposals, particularly in 2D materials-based heterostructures and quantum information science related arenas. For example, charged impurities in h-BN likely represent the limiting source of disorder in graphene/h-BN heterostructures and are detrimental to any attempts at manipulation of charged quasiparticles for quantum information science applications [48, 49, 50].
@@
From a practical perspective, these transition metal fluxes represent a substantive synthetic advance due to their ability to dissolve and recrystallize high-quality h-BN at atmospheric pressure. This greatly reduces the technical difficulty of producing single crystals via the CBF technique, enabling production without the highly specialized equipment needed for the Ba3BN2 flux [38]. An as yet unrealized benefit of these fluxes is the possibility of applying the traveling-solvent floating-zone (TSFZ) method to growth of h-BN crystals. This is particularly exciting because the nature of the TSFZ method makes it well-suited to overcoming the main synthetic challenge associated with fully realizing the technological possibilities of h-BN: producing high-volume crystals with a low defect density. As a form of continuous, directional crystallization, the TSFZ method is a route towards larger, bulk, single-crystals than what is typically produced by the CBF method – which are generally on the order of 1 mm in lateral dimensions and up to 100 $\mu$m thick – while retaining the high quality. For one, TSFZ functions intrinsically with very high temperature gradients and experiments with CBF-growth under a temperature gradient have demonstrated improved structural quality as compared to slow cooling alone [43]. Additionally, the TSFZ method has a number of benefits with respect to sample purity which can help to reduce the extrinsic contamination seen in CBF-grown crystals via e.g. cathodoluminescence and secondary ion mass spectrometry [51]. First, TSFZ is a “crucible-free” growth technique wherein the liquid flux does not come into contact with any additional material besides the one being grown. Second, TSFZ benefits from the zone-refining phenomenon, wherein impurities are retained in the molten zone due to a higher liquid solubility than in the crystallized solid [52].
@@
In this work, we report the feasibility of the TSFZ technique for crystal growth of h-BN. The effects of a variety of TSFZ growth parameters were studied, guided by efficient first-pass characterization via X-ray tomography. Two variations of TSFZ proved to be feasible for growing h-BN crystals and the grown boules range in size, with diameters of 3 — 5 mm and lengths from 2 – 10 mm. Tomography data indicates the presence of grains of variable size within the boules, with the largest observed having areas of $\approx$ 1 × 2 mm and thickness $\approx$ 0.5 mm. For optimized growth conditions (TSFZ geometry, Fe flux, $P$ = 7 bar, 0.1 mm/h growth rate) average Raman linewidths of 7.7(2) cm-1 and 1.0(1) cm-1 were obtained for the E2g modes at 1365.46(4) cm-1 and 51.78(9) cm-1, respectively. The corresponding PL spectrum shows sharp phonon-assisted free exciton peaks as well as minimal signal in the energy range corresponding to C defects ($E$ = 3.9 – 4.1 eV).
@@ 2 Crystal growth methodology
A majority of the TSFZ growths in this work were conducted using the tilting, laser-based floating-zone system at the Platform for the Accelerated Realization, Analysis and Discovery of Interface Materials (PARADIM) crystal growth facility (Crystal Systems, Inc., FD-FZ-5-200-VPO-PC), which utilizes 5 $\times$ 200 W lasers ($\lambda$ = 976 nm). Additional growths were conducted at the University of California, Santa Barbara (UCSB) using a custom-built, high-pressure, laser-based floating-zone system [53, 54] which utilizes 7 $\times$ 200 W lasers ($\lambda$ = 1070 nm). Two different growth geometries were tested to find the optimum conditions for h-BN. In standard TSFZ, a pellet of the flux is balanced on top of a seed rod made of polycrystalline h-BN, with a feed rod installed above. To initiate the growth the flux is melted, fusing it to the seed rod, and the feed rod is translated downwards until it connects to the molten flux and forms a stable molten zone. By translating the feed and seed rods downwards in tandem the feed rod dissolves into the flux and h-BN recrystallizes onto the seed, causing the molten flux to "travel". An alternative geometry – similar to Czochralski growth – is to place the feed on the bottom, melt the flux in the same manner as in TSFZ, and then dip a seed rod into the molten flux from above. Both rods are then translated upwards to simultaneously dissolve the feed rod and recrystallize h-BN. We refer to this geometry as "laser pedestal growth" (LPG) [55, 56] to distinguish it from the standard geometry, which is referred to as TSFZ. To summarize, in the TSFZ technique the feed rate (vfeed) corresponds to vU in Fig. 1, while the crystal growth rate (vgrow) corresponds to vL. In LPG the situation is reversed.
The main parameters relevant to these geometries are illustrated in Fig. 1 and include the rod diameter ($D$), rod movement rate and direction (v), rod composition, flux composition, and total gas pressure ($P$).
@@ 2.1 Flux preparation
Two fluxes with different transition metal stoichiometries were tested: Fe and FeCr. Additional variants of these fluxes were tested by adding B prior to the growth to form nominal compositions of FeB0.2 and FeCrB0.2. The fluxes were created by arc melting stoichiometric amounts of Fe (Kurt J. Lesker Co., 99.95%), Cr (Kurt J. Lesker Co., 99.95%), and amorphous B (Thermo Scientific, 98%) together under an Ar atmosphere. The Fe:B ratio was chosen based on the solubility of B in liquid Fe [57]. The resulting ingots were then sectioned into smaller amounts and arc melted again on a flat hearth to create flat-bottomed, dome-shaped pellets which were more easily balanced on the seed rod. The mass of the flux pellet for a growth was chosen based on the diameter of the seed rod: typically $\approx$ 0.2 g for a 3 mm diameter and $\approx$ 0.5 g for a 5 - 6 mm diameter. These masses yielded a sufficient volume of molten flux such that the full surface of the rods could be fused to the flux.
@@ 2.2 Gas environment
In the growths conducted at PARADIM, concurrent flows of N2 and Ar:H2 = 95:5 (by volume) were utilized. The N2 was supplied either from a standard gas cylinder (Airgas, 99.999%) or a high-pressure liquid dewar (Airgas, 99.999%). In later growths, the N2 was passed through a room temperature getter (Nupure, Eliminator Model 200 CA). Growths were conducted at a total pressure $P$ = 1, 5 or 7 bar using a N2 flow rate of 0.5 L/min and an Ar:H2 flow rate of 0.05 L/min, yielding a N2 partial pressure of $P_{N2}\approx 0.9P$. The chamber was purged at these flow rates for 15 minutes prior to setting the final pressure. For the growths at UCSB a static pressure of pure N2 was utilized because the Inconel pressure chamber is not compatible with H2. Growths were conducted at $P=P_{N2}$ = 50 bar with the N2 supplied from a standard gas cylinder (Airgas, 99.999%) and flowed through a room temperature getter (Nupure, Eliminator Model 1000 CG). The chamber was initially purged at atmospheric pressure using a 1 L/min flow of N2 for 15 minutes; the pressure was then cycled several times between 30 and 3 bar before setting the final pressure.
@@ 2.3 Rod composition, size, and movement rate
When h-BN was used as the feed rod, commercially available hot-pressed, polycrystalline samples were utilized. For the growths at PARADIM the rods (Momentive, HBN-type) contained roughly 5 weight % B2O3, as received. Prior to growth, these polycrystalline h-BN rods were annealed at 1400 °C under vacuum (5 $\times$ 10-5 torr) for 24 hours, resulting in significant evaporation of B2O3. SEM/EDS analysis of these rods post degassing showed no sign of any elements besides B and N. TSFZ growths were performed with both 3 mm and 6 mm diameter polycrystalline h-BN feed and seed rods. In some cases, a pure B crystal (diameter 2 – 5 mm) was utilized as the feed rod (see below). For the LPG geometry growths, the feed was either a 5 mm polycrystalline h-BN rod or a 5 – 6 mm B crystal. These growths were seeded using a boule from a previous TSFZ growth (either the sample shown in Fig. 2c or that in Fig. S1a) which was kept attached to its polycrystalline seed rod. For the TSFZ growths at UCSB, 3 mm diameter polycrystalline h-BN rods (Premier Ceramics) were used as the feed and seed. These rods were quoted by the supplier as “binder free” and were not degassed before use. For the UCSB and PARADIM growths with polycrystalline h-BN the tip of the feed rod was sanded into a thin conical shape prior to use. Additionally, a small divot was drilled into the polycrystalline seed rod to secure the flux prior to melting.
@@
In all PARADIM TSFZ h-BN growths, vgrow was set to 0.1 mm/h and vfeed was set to 0.1 – 0.4 mm/h depending on the ratio between the feed and seed rod diameters and the composition of the feed rod. In LPG growths, vgrow was set to 0.1 mm/h. The feed rod was not moved during LPG growth because the diameter of the seed rod was substantially smaller than that of the feed. For the UCSB growths, vgrow = vfeed = 0.4 mm/h was used since this was the slowest speed achievable by the translation system. In both PARADIM and UCSB growths, seed rotation rates were between 4 – 8 rpm and feed rotation rates were between 8 – 15 rpm, depending on the status of the growth in progress.

### v2

@@ Abstract
Large, high-purity single-crystals of hexagonal BN (h-BN) are essential for exploiting its many desirable and interesting properties. Here, we demonstrate via X-ray tomography, X-ray diffraction and scanning electron microscopy that h-BN crystals can be grown by traveling-solvent floating-zone (TSFZ). The diameters of grown boules range from 3 – 5 mm with lengths from 2 – 10 mm. Tomography indicates variable grain sizes within the boules, with the largest having areas of $\approx$ 1 mm $\times$ 2 mm and thickness $\approx$ 0.5 mm. Although the boules contain macroscale flux inclusions, the h-BN lattice itself is of high quality for samples grown under optimized conditions. The currently optimized growth procedure employs an Fe flux, moderate N2 pressure ($P_{N2}\approx$ 6 bar), and a growth rate of 0.1 mm/h. Raman spectroscopy for an optimized sample gives an average linewidth of 7.7(2) cm-1 for the E2g intralayer mode at 1365.46(4) cm-1 and 1.0(1) cm-1 for the E2g interlayer shear mode at 51.78(9) cm-1. The corresponding photoluminescence spectrum shows sharp phonon-assisted free exciton peaks and minimal signal in the energy range corresponding to carbon-related defects ($E$ = 3.9 – 4.1 eV). Our work demonstrates the viability of growing h-BN by the TSFZ technique, thereby opening a new route towards larger, high-quality crystals and advancing the state of h-BN related research.
@@ 1 Introduction
The utility of hexagonal BN (h-BN) has been recognized since at least the 1950’s, particularly for high-temperature applications, where its chemical inertness, thermodynamic stability and high electrical resistivity are extremely useful. The arrival of graphene [1, 2] subsequently stimulated additional interest in h-BN as an ideal insulating substrate for device applications [3]. h-BN is an excellent substrate for graphene-based devices due to the combination of a small lattice mismatch and atomic-level flatness (i.e. no dangling bonds) with good dielectric [1, 2, 4] and thermal transport [5, 6] properties. The chemical inertness and oxidation resistance of h-BN [7] also make it valuable for capping and encapsulation purposes [8]. Furthermore, h-BN can also take an active role, as in graphene/h-BN moiré superlattices [9, 10, 11, 12]. The same properties which make h-BN ideal for graphene also make it a useful platform for device creation using other two-dimensional (2D) materials and it has seen widespread deployment in this context.
More recently, h-BN is under study for a host of additional applications, including, among others, photonics, flexible electronics, non-linear optics and quantum information science and sensing technologies [13, 14, 15, 16]. A number of these applications exploit the intrinsic properties of h-BN on their own. As a naturally hyperbolic  [^1] Here, hyperbolic indicates that the relative permittivity changes sign as well as magnitude along different crystallographic directions. birefringent material in the mid-infrared range h-BN is well-suited to many photonics applications [17], such as confinement and manipulation of phonon polaritons [18]. Relatedly, point defects within h-BN have also begun to generate significant interest as single photon emitters [19, 20, 21], which have the potential to serve as an optical platform for quantum information [22] and as a spin defect for use in atomic-resolution quantum sensing [23, 24, 25].
@@
Accordingly, the primary source of high-quality h-BN single crystals for the most demanding experiments has principally been the crucible-based flux (CBF) method developed by Taniguchi and Watanabe [38], which utilizes a Ba3BN2 flux in Mo crucibles at high temperature and GPa-scale pressures (commonly referred to as the high-pressure, high-temperature, or HPHT, technique). However, subsequent work has demonstrated that single-crystal h-BN samples of comparable quality and size can be made with various transition metal based fluxes such Fe-Cr or Ni-Cr alloys [39, 40, 41, 42, 43, 44, 45] (referred to as the atmospheric-pressure, high temperature, or APHT, technique). These crystals have been utilized successfully in, among others, a number of interesting photonics experiments [46, 47]. Despite these accomplishments, it is clear that further improvements are needed to enable the most ambitious technological proposals, particularly in 2D materials-based heterostructures and quantum information science related arenas. For example, charged impurities in h-BN likely represent the limiting source of disorder in graphene/h-BN heterostructures and are detrimental to any attempts at manipulation of charged quasiparticles for quantum information science applications [48, 49, 50].
@@
From a practical perspective, these transition metal fluxes represent a substantive synthetic advance due to their ability to dissolve and recrystallize high-quality h-BN at atmospheric pressure. This greatly reduces the technical difficulty of producing single crystals via the CBF technique, enabling production without the highly specialized equipment needed for the Ba3BN2 flux [38]. An as yet unrealized benefit of these fluxes is the possibility of applying the traveling-solvent floating-zone (TSFZ) method to growth of h-BN crystals. This is particularly exciting because the nature of the TSFZ method makes it well-suited to overcoming the main synthetic challenge associated with fully realizing the technological possibilities of h-BN: producing high-volume crystals with a low defect density. As a form of continuous, directional crystallization, the TSFZ method is a route towards larger, bulk, single-crystals than what is typically produced by the CBF method – which are generally on the order of 1 mm in lateral dimensions and up to 100 $\mu$m thick – while retaining the high quality. For one, TSFZ functions intrinsically with very high temperature gradients and experiments with CBF-growth under a temperature gradient have demonstrated improved structural quality as compared to slow cooling alone [43].
The TSFZ method also has a number of benefits with respect to sample purity which can help to reduce the extrinsic contamination seen in CBF-grown crystals via e.g. cathodoluminescence and secondary ion mass spectrometry [51]. First, TSFZ is a "crucible-free" growth technique in the sense that the liquid flux does not come into contact with any additional material besides the one being grown. Second, TSFZ benefits from the zone-refining phenomenon, wherein impurities are retained in the molten zone due to a higher liquid solubility than in the rcrystallized solid [52].
@@
In this work, we report the feasibility of the TSFZ technique for crystal growth of h-BN. The effects of a variety of TSFZ growth parameters were studied, guided by efficient first-pass characterization via X-ray tomography. Two variations of TSFZ proved to be feasible for growing h-BN crystals and the grown boules range in size, with diameters of 3 – 5 mm and lengths from 2 – 10 mm. Tomography data indicates the presence of grains of variable size within the boules, with the largest observed having an area of $\approx$ 1 × 2 mm and thickness $\approx$ 0.5 mm. For optimized growth conditions (TSFZ geometry, Fe flux, $P$ = 7 bar, 0.1 mm/h growth rate) average Raman linewidths of 7.7(2) cm-1 and 1.0(1) cm-1 were obtained for the E2g modes at 1365.46(4) cm-1 and 51.78(9) cm-1, respectively. The corresponding PL spectrum shows sharp phonon-assisted free exciton peaks as well as minimal signal in the energy range corresponding to carbon defects ($E$ = 3.9 – 4.1 eV).
@@ 2 Crystal growth methodology
A majority of the TSFZ growths in this work were conducted using the tilting, laser-based floating-zone system at the Platform for the Accelerated Realization, Analysis and Discovery of Interface Materials (PARADIM) crystal growth facility (Crystal Systems, Inc., FD-FZ-5-200-VPO-PC), which utilizes 5 $\times$ 200 W lasers ($\lambda$ = 976 nm). Additional growths were conducted at the University of California, Santa Barbara (UCSB) using a custom-built, high-pressure, laser-based floating-zone system [53, 54] which utilizes 7 $\times$ 200 W lasers ($\lambda$ = 1070 nm). Two different growth geometries were tested to find the optimum conditions for h-BN. In standard TSFZ, a pellet of the flux is balanced on top of a seed rod made of polycrystalline h-BN, with a feed rod installed above. To initiate the growth the flux is melted, fusing it to the seed rod, and the feed rod is translated downwards until it connects to the molten flux and forms a stable molten zone. By translating the feed and seed rods downwards in tandem the feed rod dissolves into the flux and h-BN recrystallizes onto the seed, causing the molten flux to "travel". An alternative geometry – similar to Czochralski growth – is to place the feed on the bottom, melt the flux in the same manner as in TSFZ, and then dip a seed rod into the molten flux from above. The seed rod is then translated upwards to recrystallize h-BN from the zone while the feed rod may or may not be translated depending on its size relative to the seed. We refer to this geometry as "laser pedestal growth" (LPG) [55, 56] to distinguish it from the standard geometry, which is referred to as TSFZ. To summarize, in the TSFZ technique the feed rate (vfeed) corresponds to vU in Fig. 1, while the crystal growth rate (vgrow) corresponds to vL. In LPG the situation is reversed.
The main parameters relevant to these geometries are illustrated in Fig. 1 and include the rod diameter ($D$), rod movement rate and direction (v), rod composition, flux composition, and total gas pressure ($P$). In Table S1 in the Supplemental Information (SI) we show a summary of the growth parameters for all samples considered in this work, which are referred to by their sample numbers (e.g. 1, 2, 3) in the following.
@@ 2.1 Flux preparation
Two fluxes with different transition metal stoichiometries were tested: Fe and FeCr. Additional variants of these fluxes were tested by adding B prior to the growth (presaturation) to form nominal compositions of FeB0.2 and FeCrB0.2. The fluxes were created by arc melting stoichiometric amounts of Fe (Kurt J. Lesker Co., 99.95%), Cr (Kurt J. Lesker Co., 99.95%), and amorphous B (Thermo Scientific, 98%) together under an Ar atmosphere. The Fe:B ratio was chosen based on the solubility of B in liquid Fe [57]. The resulting ingots were then sectioned into smaller amounts and arc melted again on a flat hearth to create flat-bottomed, dome-shaped pellets which were more easily balanced on the seed rod. The mass of the flux pellet for a growth was chosen based on the diameter of the seed rod: typically $\approx$ 0.2 g for a 3 mm diameter and $\approx$ 0.5 g for a 5 — 6 mm diameter. These masses yielded a sufficient volume of molten flux such that the full surface of the rods could be fused to the flux.
@@ 2.2 Gas environment
In the growths conducted at PARADIM, concurrent flows of N2 and Ar:H2 = 95:5 (by volume) were utilized. The N2 was supplied either from a standard gas cylinder (Airgas, 99.999%) or a high-pressure liquid dewar (Airgas, 99.999%). In later growths, the N2 was passed through a room temperature getter (Nupure, Eliminator Model 200 CA). Growths were conducted at a total pressure $P$ = 1, 5 or 7 bar using a N2 flow rate of 0.5 L/min and an Ar:H2 flow rate of 0.05 L/min, yielding a N2 partial pressure of $P_{N2}\approx 0.9P$. Note that in the following we always refer to the total pressure used in a particular growth. The chamber was purged at these flow rates for 15 minutes prior to setting the final pressure. For the growths at UCSB a static pressure of pure N2 was utilized because the Inconel pressure chamber is not compatible with H2. Growths were conducted at $P=P_{N2}$ = 50 bar with the N2 supplied from a standard gas cylinder (Airgas, 99.999%) and flowed through a room temperature getter (Nupure, Eliminator Model 1000 CG). The chamber was initially purged at atmospheric pressure using a 1 L/min flow of N2 for 15 minutes; the pressure was then cycled several times between 30 and 3 bar before setting the final pressure.
@@ 2.3 Rod composition, size, and movement rate
When h-BN was used as the feed rod, commercially available hot-pressed, polycrystalline samples were utilized. For the growths at PARADIM the rods (Momentive, HBN-type) contained roughly 5 weight % B2O3, as received. Prior to growth, these polycrystalline h-BN rods were annealed at 1400 °C under vacuum (5 $\times$ 10-5 torr) for 24 hours, resulting in significant evaporation of B2O3. SEM/EDS analysis of these rods post degassing showed no sign of any elements besides B and N. TSFZ growths were performed with both 3 mm and 6 mm diameter polycrystalline h-BN feed and seed rods. In some cases, a pure B crystal (diameter 2 – 5 mm) was utilized as the feed rod (see below). For the LPG geometry growths, the feed was either a 5 mm polycrystalline h-BN rod or a 5 – 6 mm B crystal. These growths were seeded using a boule from a previous TSFZ growth (either the sample shown in Fig. 2b or that in Fig. S1a) which was kept attached to its polycrystalline seed rod. For the TSFZ growths at UCSB, 3 mm diameter polycrystalline h-BN rods (Premier Ceramics) were used as the feed and seed. These rods were quoted by the supplier as “binder free” and were not degassed before use. For the UCSB and PARADIM TSFZ growths with polycrystalline h-BN rods the tip of the feed rod was sanded into a thin conical shape prior to use. Additionally, a small divot was drilled into the polycrystalline seed rod to secure the flux prior to melting.
@@
In all PARADIM TSFZ h-BN growths, vgrow was set to 0.1 mm/h and vfeed was set to 0.1 – 0.4 mm/h depending on the ratio between the feed and seed rod diameters and the composition of the feed rod. In LPG growths, vgrow was set to 0.1 mm/h. The feed rod was not moved during LPG growth because the diameter of the seed rod was substantially smaller than that of the feed. For the growths at UCSB it was not possible to use the same vgrow = 0.1 mm/h because the slowest translation speed achievable with the high-pressure compatible drives is 0.4 mm/h. As a result, the high-pressure growth (sample 3) utilized vgrow = vfeed = 0.4 mm/h .
In the TSFZ growths at PARADIM and UCSB, the seed rotation rate was 8 rpm and the feed rotation rate was typically 9 rpm. Over the course of growths with a h-BN feed the interface between the feed rod and flux occasionally deteriorated due to uneven dissolution of the feed. A faster feed rotation rate of 15 rpm was then briefly employed ($\approx$ 1 h) until the interface quality improved. For the LPG growths at PARADIM the rotation rates were as follows: sample 6 used a feed rotation rate of 6 rpm and a seed rotation rate of 8 rpm. Sample 10 used a feed rotation of 4 rpm and a seed rotation rate of 12 rpm.